# Multidata Causal Discovery for Statistical Hurricane Intensity Forecasting

### v2

@@ Abstract
Improving statistical forecasts of Tropical Cyclone (TC) intensity is limited by complex nonlinear interactions and difficulty in identifying relevant predictors. Conventional methods prioritize correlation or fit, often overlooking confounding variables and limiting generalizability to unseen TCs. To address this, we leverage a multidata causal discovery framework with a replicated dataset based on Statistical Hurricane Intensity Prediction Scheme (SHIPS) using ERA5 meteorological reanalysis. We conduct multiple experiments to identify and select predictors causally linked to TC intensity changes. We then train multiple linear regression models to compare causal feature selection with no selection, correlation, and random forest feature importance across five forecast lead times from 1 to 5 days (24–120 hours). Causal feature selection consistently outperforms on unseen test cases, especially for lead times shorter than 3 days. The causal features primarily include vertical shear, mid-tropospheric potential vorticity and surface moisture conditions, which are physically significant yet often underutilized in TC intensity predictions. We build an extended predictor set (SHIPS++) by adding selected features to the standard SHIPS predictors. SHIPS++ yields increased short-term predictive skill at lead times of 24, 48, and 72 hours. Adding nonlinearity using multilayer perceptron further extends skill to longer lead times, despite our framework being purely regional and not requiring global forecast data. Operational SHIPS tests confirm that three of the six added causally discovered predictors improve forecast skill, with the largest gains at longer lead times. Our results demonstrate that causal discovery improves TC intensity prediction and pave the way toward more empirical forecasts.
@@
: As data-driven TC forecasting tools become more sophisticated, it is critical to ensure that they make accurate forecasts for the right reasons, especially for tropical storms not included during training. Here, we use a technique called multidata causal discovery, which identifies a unique set of predictors driving TC intensity changes over the next five days by analyzing multiple realizations of the same underlying phenomenon: tropical storm in the North Atlantic. Adding these predictors to an operationally used statistical TC intensity prediction scheme improves forecasting accuracy. This shows that selecting predictors based on cause-and-effect, not just correlation, can result in improved forecasts that are more generalizable as well as interpretable, since the causally discovered predictors are physically consistent.
@@ 1 Introduction
Extensive research on tropical cyclones (TCs) has advanced our understanding of key atmospheric and oceanic processes that control cyclogenesis and intensification Gray (1998); Elsberry (1995, 2014). However, the continued rise in coastal populations and the increasing risks of wind gusts, storm surge, extreme rainfall, and severe weather Pielke Jr and Landsea (1998); Crossett et al. (2004) underscore the need for more accurate and robust TC intensity forecasts. Despite breakthroughs in short-range to sub-seasonal prediction Camargo et al. (2019); Heming et al. (2019); Tang et al. (2020), accurately predicting rapid storm intensity changes remains challenging DeMaria and Kaplan (1999); Wang and Wu (2004); DeMaria et al. (2021) especially beyond 24 hours, due to error growth in initial conditions, incomplete representation of key physical processes, and limitations in data assimilation Emanuel and Zhang (2016). For example, from 2010 - 2019, the National Hurricane Center reported mean TC intensity errors ranging from 5.7 knots ($\sim$ 2.6 m/s) for 12 hour to 12 knots ($\sim$ 6 m/s) for 48 hour forecasts Cangialosi et al. (2020). In addition to numerical weather prediction systems, statistical models play a significant role in TC forecasts by improving the skill of multi-model ensemble-based consensus forecasts DeMaria et al. (2022). Both statistical and dynamical models struggle to estimate the rapid intensification (RI) as multiscale air-sea interactions and radiative-convective feedbacks leading to rapid intensity changes are not fully understood Wang (2012). RI events account for 20 - 30% of the total intensity forecast error variance Trabing and Bell (2020), and peak intensity of rapidly intensifying TCs is routinely underestimated Torn and DeMaria (2021). Although statistical schemes outperform high-resolution dynamical models in probabilistic RI prediction Torn and DeMaria (2021), dynamical model intensity forecasts, including during RI events, improve substantially when advanced configurations such as storm-following nests are employed Alaka Jr et al. (2022).
@@
Statistical Hurricane Intensity Prediction Scheme (SHIPS) integrates large-scale predictors from climatology, persistence, and synoptic predictors to estimate TC intensity DeMaria and Kaplan (1994); Marks and DeMaria (2003) and its forecast skill has gradually improved over time. SHIPS evolved from a “statistical-synoptic” to a “statistical-dynamical” framework through the incorporation of synoptic environmental conditions from dynamical models (e.g. GFS) DeMaria et al. (2022). The combined consensus forecast exhibits the best skill for TC intensity prediction, with SHIPS adding value to operational intensity prediction efforts Cangialosi et al. (2020). SHIPS includes an operational probabilistic RI index based on large-scale environmental predictors Kaplan et al. (2010, 2015). Machine learning approaches have further improved 24 hour RI prediction skill, with SHIPS-compatible extensions using satellite-derived predictors, hybrid deep learning–gradient boosting frameworks, and operational consensus models Su et al. (2020); Boussioux et al. (2022); Ko et al. (2023).
@@
In this study, we address this limitation by adopting a causal discovery framework Runge (2018); Runge et al. (2019a, b); Gerhardus and Runge (2022); Runge et al. (2023) to objectively identify new environmental predictors for inclusion in the operational SHIPS developmental dataset. In causal discovery, relationships among variables are represented using a causal graph, where nodes correspond to physical variables, and directed edges denote direct causal influences, potentially with time lags. While causal discovery aims to infer the full causal structure among a set of variables, causal feature selection instead uses a causal discovery algorithm to identify a subset of physically meaningful predictors that are most relevant for predicting a target variable, without reconstructing the full causal graph. Climate science studies have applied linear and nonlinear causal discovery methods Ebert-Uphoff and Deng (2012); Runge et al. (2014b, 2019b) to identify spurious associations in statistical prediction models arising from common drivers or indirect associations. Once the causal graph is known, the strength of the links between the physical variables can be determined with causal inference Ricard et al. (2025), complementing data-driven techniques Runge et al. (2023). Despite these advances, causality is still commonly inferred using lagged correlation analysis, which does not resolve the directionality of the relationships between variables McGraw and Barnes (2018). As such, lagged correlation methods are susceptible to non-causal correlations from autocorrelation effects, indirect connections through a third process, or a shared driver hindering interpretability in statistical prediction models Runge et al. (2014a); Barnes and Screen (2015). Understanding the directional causality between atmospheric variables hence requires additional methodologies or causal inference approaches beyond the scope of lagged correlation analyses. Several studies have demonstrated the potential of causal discovery methods in tropical meteorology. For example, Wijnands et al. (2016) used the graphical model-based structure learning approach of the Peter-Clark (PC) causal discovery algorithm to identify key variables for tropical cyclogenesis prediction and found that predictive modeling using logistic regression employing the highest ranked variables improved statistical predictive skills. Additionally, the PCMCI+ causal algorithm — which combines the PC algorithm with Momentary Conditional Independence test Runge (2020) — has been used to identify precursor regions useful for improving seasonal TC frequency forecasts Latos et al. (2024). Motivated by these advances, we apply the PC algorithm within a multidata causal discovery framework to identify statistically significant causal links and select predictors with direct causal influence on TC intensity change for the operational SHIPS model. This approach treats individual Atlantic tropical cyclones as multiple realizations of the same underlying physical process, assuming that their intensity evolution shares a common causal structure that can be represented by a single causal graph. The multidata PC algorithm requires multivariate time series of all candidate predictors for each storm, motivating our replication of the SHIPS developmental dataset using meteorological analyses, as described in Section 2. After detailing the causal methodology and experimental design in Section 3, we demonstrate in Sections 44.1 and 44.2 that causal feature selection outperforms other feature selection methods, particularly in short-range time-scales. These improvements are robust in both reanalysis-based experiments (Section 44.3), where we also show that the causal predictors nonlinearly drive intensity (Section 44.4), and in operational-like settings using real-time global model analysis and forecast fields (Section 44.6). In Section 44.5, we present a case study of Hurricane Larry to illustrate the physical relevance of the causal links between the selected predictors and short-term TC intensification.
@@ 2 Data Sources and Preprocessing
Our analysis is based on two complementary experimental designs: (i) reanalysis-based experiments using the high-resolution ECMWF Reanalysis Version 5 (ERA5) Hersbach et al. (2018, 2020), and (ii) experiments using SHIPS developmental to test the added value of causal predictors compared to the operational SHIPS predictors. To define TC-centered domains and intensity change targets, we rely on the International Best Track Archive for Climate Stewardship (IBTrACS), which provides reference track and intensity data for each TC.
@@ 2.2 ERA5 and TC PRIMED Datasets
Causal discovery is restricted to SHIPS predictors that can be replicated using ERA5; satellite-only predictors such as GOES brightness temperature and ocean heat content are excluded to ensure consistency. The ERA5 predictor inputs consist of six-hourly, area-averaged time series extracted from TC-following domains centered on IBTrACS best-track positions. Predictors are computed for multiple radial regions around the storm center to capture inner core and outer area conditions: inner-core variables are averaged over 0–2 (approximately 0–200 km), while outer-area predictors are averaged from 200–800 km or up to 1000 km, following conventions used in the extended SHIPS developmental dataset. To avoid backward-in-time causal dependencies, we only consider predictors at analysis (00 hour) time unlike operational SHIPS setup, which incorporates dynamical model forecast fields. As a result, forecast skill at longer lead times may be underestimated when synoptic-scale evolution is important. To expand the predictor set further, we use the Tropical Cyclone Precipitation, Infrared, Microwave, and Environmental Dataset (TC PRIMED) Razin and others (2023). TC PRIMED uses ERA5 to reconstruct operational SHIPS predictors as well as environmental and thermodynamic variables across multiple pressure levels and radial regions. An important limitation of the area-averaging approach is that it does not capture azimuthal asymmetries, which may be relevant for sheared or rapidly intensifying TCs Hazelton et al. (2017); Rogers et al. (2013); McNeely et al. (2020) and could be considered in future experiments.
@@
A key requirement for applying causal discovery to time series is the assumption of causal stationarity, i.e., the causal relationships between variables remain invariant over time within a given TC and across multiple storms in the multi-data setting. This allows pooling of statistical evidence across time and the full training set for conditional independence testing Ganesh et al. (2023). Accordingly, TC time series are aligned relative to the time of minimum central pressure to satisfy the causal stationarity assumption. The mean sea level pressure (MSLP) time series for each storm in the training set are smoothed using a Gaussian filter ($\sigma=3\times 6\mathrm{hr}$) to reduce noise and the time of minimum central pressure is used as the reference point for alignment. For the shorter time series, we append NaN values on either sides to make sure that the lengths are consistent with the longest time series with minimum pressure in the middle so that the evolution of different storms could be compared on a common reference frame. All predictors are subsequently standardized using the mean and standard deviation computed from the training set. The final ERA5 dataset includes 214 predictors and the intensity-change target (DELV) at each forecast lead time. A total of 216 TCs from 2000–2019 are used within a storm-based cross-validation framework with predefined, non-overlapping training and validation splits, while an independent test set comprises 31 TCs from 2020–2021 including Hurricane Wilma (2005), retained as a particularly intense and challenging case.
@@ 2.3 SHIPS Developmental Dataset
To evaluate how causally informed predictors improve operational forecasting, we extract SHIPS predictors (Table S1) for the same 247 North Atlantic TC cases (2000-2021) used in the ERA5 experiments. We only used predictor values from the 00 hour initialization time in the SHIPS developmental dataset derived from GFS analysis fields, which represent the atmospheric conditions (i.e., observations assimilated into the GFS model) at forecast initialization time. This setup enables direct analogy between the SHIPS developmental dataset and the ERA5-derived dataset used for causal predictor discovery, as both rely on analysis time atmospheric fields within a purely statistical framework. We compare the generalization skill of statistical models trained using the original 21 SHIPS predictors (“original SHIPS”) with models trained on the same predictors augmented with causally selected predictors (“SHIPS+”). These additional predictors are identified through ERA5-based causal discovery experiments but are computed from GFS analysis fields in exactly the same way as the original SHIPS predictors. This setup tests whether the predictors selected in the ERA5 replication setting improve forecast skill when integrated into the operational SHIPS framework.
@@ 3.1 Causal Feature Selection
Causal feature selection is performed using the open-source TiGraMITe causal discovery framework Runge (2018); Runge et al. (2019b) which has been widely used to discover causal relationships from climate data Kretschmer et al. (2016, 2018); Samarasinghe et al. (2021). We employ the multidata PC (M-PC) algorithm, a modified version of the first step of the PCMCI algorithm Runge et al. (2019b); Runge (2020), to discover predictors of TC intensity change while eliminating spurious statistical associations. Unlike PCMCI+, which aims to infer the full causal graph including contemporaneous causal links, this work uses only the PC algorithm for causal feature selection Iglesias-Suarez et al. (2024); Runge et al. (2015). The “M” in M-PC indicates the multidata approach, where an ensemble of time series from multiple TCs is analyzed to identify a single set of causal predictors. Non-causal drivers within the ERA5 predictor set are detected and excluded through conditional independence test using linear partial correlation Ganesh et al. (2023). In our setup, this is done without considering time-lagged variables — each predictor is evaluated at a single time-step for the given forecast lead time. By filtering out non-causal predictors in the statistical models, this approach aims to improve the generalizability of statistical models to unseen TC cases.

### v3

@@ Abstract
Improving statistical forecasts of tropical cyclone (TC) intensity is limited by complex nonlinear interactions and difficulty in identifying relevant predictors. Conventional methods prioritize correlation or fit, often overlooking confounding variables and limiting generalizability to unseen TCs. To address this, we leverage a multidata causal discovery framework with a replicated dataset based on Statistical Hurricane Intensity Prediction Scheme (SHIPS) using ERA5 meteorological reanalysis. We conduct experiments to identify and select predictors causally linked to TC intensity changes. We then train multiple linear regression models to compare causal feature selection with correlation, random forest feature importance, and no feature selection, across five forecast lead times from 1 to 5 days (24–120 hours). Causal feature selection consistently outperforms on unseen test cases, especially for lead times shorter than 3 days. Top causal features include vertical shear, mid-tropospheric potential vorticity and surface moisture conditions, which are physically significant yet often underutilized in TC intensity predictions. We build an extended predictor set (SHIPS++) by adding selected features to the standard SHIPS predictors. SHIPS++ yields increased short-term predictive skill at lead times of 24, 48, and 72 hours. Adding nonlinearity using a multilayer perceptron further extends skill to longer lead times, despite our framework being purely regional and not requiring global forecast data. Operational SHIPS tests confirm that three of the six added causally discovered predictors improve forecast skill, with the largest gains at longer lead times. Our results demonstrate that causal discovery improves TC intensity prediction and pave the way toward more empirical forecasts.
@@
: As data-driven tropical cyclone (TC) forecasting tools become more sophisticated, it is critical to ensure that they make accurate forecasts for the right reasons, especially for TCs not included during training. Here, we use a technique called multidata causal discovery, which identifies a unique set of predictors driving TC intensity changes over the next five days by analyzing multiple realizations of the same underlying phenomenon: TCs in the North Atlantic. Adding these predictors to an operationally used statistical TC intensity prediction scheme improves accuracy of forecasts. This shows that selecting predictors based on cause-and-effect, not just correlation, can result in improved forecasts that are more generalizable as well as interpretable, since the causally discovered predictors are physically consistent.
@@ 1 Introduction
Extensive research on tropical cyclones (TCs) has advanced our understanding of key atmospheric and oceanic processes that control cyclogenesis and intensification Gray (1998); Elsberry (1995, 2014). However, the continued rise in coastal populations and the increasing risks of wind gusts, storm surge, extreme rainfall, and severe weather Pielke Jr and Landsea (1998); Crossett et al. (2004) underscore the need for more accurate and robust TC intensity forecasts. Despite breakthroughs in short-range to subseasonal prediction Camargo et al. (2019); Heming et al. (2019); Tang et al. (2020), accurately predicting rapid TC intensity changes remains challenging DeMaria and Kaplan (1999); Wang and Wu (2004); DeMaria et al. (2021) especially beyond 24 hours, due to error growth in initial conditions, incomplete representation of key physical processes, and limitations in data assimilation Emanuel and Zhang (2016). For example, recent National Hurricane Center verification statistics (2020–2024) indicate mean TC intensity errors of 5.1 knots ($\sim$ 2.6 m s-1) at 12-hour lead time and 10.0 knots ($\sim$ 5.1 m s-1) at 48-hour lead time National Hurricane Center (2025). In addition to numerical weather prediction systems, statistical models play a significant role in TC forecasts by improving the skill of multi-model ensemble-based consensus forecasts DeMaria et al. (2022). Both statistical and dynamical models struggle to estimate rapid intensification (RI) as multiscale air-sea interactions and radiative-convective feedbacks leading to intensity changes are not fully understood Wang (2012). RI events account for 20 - 30% of intensity forecast error variance Trabing and Bell (2020), and peak intensity of rapidly intensifying TCs is routinely underestimated Torn and DeMaria (2021). Although statistical schemes outperform high-resolution dynamical models in probabilistic RI prediction Torn and DeMaria (2021), dynamical model intensity forecasts, including during RI events, improve substantially when advanced configurations such as storm-following nests are employed Alaka Jr et al. (2022).
@@
Statistical Hurricane Intensity Prediction Scheme (SHIPS) integrates large-scale predictors from climatology, persistence, and synoptic predictors to estimate TC intensity DeMaria and Kaplan (1994); Marks and DeMaria (2003) and its forecast skill has gradually improved over time. SHIPS evolved from a “statistical-synoptic” to a “statistical-dynamical” framework through the incorporation of synoptic environmental conditions from dynamical models (e.g. GFS) DeMaria et al. (2022). The combined consensus forecast exhibits the best skill for TC intensity prediction, with SHIPS adding value to operational intensity prediction efforts Cangialosi et al. (2020). SHIPS includes an operational probabilistic RI index based on large-scale environmental predictors Kaplan et al. (2010, 2015). Machine learning approaches have further improved 24-hour RI prediction skill, with SHIPS-compatible extensions using satellite-derived predictors, hybrid deep learning–gradient boosting frameworks, and operational consensus models Su et al. (2020); Boussioux et al. (2022); Ko et al. (2023).
@@
We address this limitation by adopting a causal discovery framework Runge (2018); Runge et al. (2019a, b); Gerhardus and Runge (2022); Runge et al. (2023) to objectively identify new environmental predictors for inclusion in the operational SHIPS developmental dataset. In causal discovery, relationships among variables are represented using a causal graph, where nodes correspond to physical variables, and directed edges denote direct causal influences, potentially with time lags. While causal discovery aims to infer the full causal structure among a set of variables, causal feature selection instead uses a causal discovery algorithm to identify a subset of physically meaningful predictors that are most relevant for predicting a target variable, without reconstructing the full causal graph. Climate science studies have applied linear and nonlinear causal discovery methods Ebert-Uphoff and Deng (2012); Runge et al. (2014b, 2019b) to identify spurious associations in statistical prediction models arising from common drivers or indirect associations. Once the causal graph is known, the strength of the links between the physical variables can be determined with causal inference Ricard et al. (2025), complementing data-driven techniques Runge et al. (2023). Despite these advances, causality is still commonly inferred using lagged correlation analysis, which does not resolve the directionality of the relationships between variables McGraw and Barnes (2018). As such, lagged correlation methods are susceptible to non-causal correlations from autocorrelation effects, indirect connections through a third process, or a shared driver hindering interpretability in statistical prediction models Runge et al. (2014a); Barnes and Screen (2015). Understanding the directional causality between atmospheric variables hence requires additional methodologies or causal inference approaches beyond the scope of lagged correlation analyses. Several studies have demonstrated the potential of causal discovery methods in tropical meteorology. For example, Wijnands et al. (2016) used the graphical model-based structure learning approach of the Peter-Clark (PC) causal discovery algorithm to identify key variables for tropical cyclogenesis prediction and found that predictive modeling using logistic regression employing the highest ranked variables improved statistical predictive skills. Additionally, the PCMCI+ causal algorithm — which combines the PC algorithm with Momentary Conditional Independence test Runge (2020) — has been used to identify precursor regions useful for improving seasonal TC frequency forecasts Latos et al. (2024). Motivated by these advances, we apply the PC algorithm within a multidata causal discovery framework to identify statistically significant causal links and select predictors with direct causal influence on TC intensity change for the operational SHIPS model. This approach treats individual Atlantic TCs as multiple realizations of the same underlying physical process, assuming that their intensity evolution shares a common causal structure that can be represented by a single causal graph. The multidata PC algorithm requires multivariate time series of all candidate predictors for each storm, motivating our replication of the SHIPS developmental dataset using meteorological analyses, as described in Section 2. After detailing the causal methodology and experimental design in Section 3, we demonstrate in Sections 44.1 and 44.2 that causal feature selection outperforms other feature selection methods, particularly in short-range time-scales. These improvements are robust in both reanalysis-based experiments (Section 44.3), where we also show that the causal predictors nonlinearly drive intensity (Section 44.4), and in operational-like settings using real-time global model analysis and forecast fields (Section 44.6). In Section 44.5, we present a case study of Hurricane Larry to illustrate the physical relevance of the causal links between the selected predictors and short-term TC intensification.
@@ 2 Data Sources and Preprocessing
Our analysis is based on two complementary experimental designs: (i) reanalysis-based experiments using the high-resolution ECMWF Reanalysis Version 5 (ERA5) Hersbach et al. (2018, 2020), and (ii) experiments using SHIPS developmental dataset to test the added value of causal predictors compared to the operational SHIPS predictors. To define TC-centered domains and intensity change targets, we rely on the International Best Track Archive for Climate Stewardship (IBTrACS), which provides reference track and intensity data for each TC.
@@ 2.2 ERA5 and TC PRIMED Datasets
Causal discovery is restricted to SHIPS predictors that can be replicated using ERA5; satellite-only predictors such as GOES brightness temperature and ocean heat content are excluded to ensure consistency. The ERA5 predictor inputs consist of six-hourly, area-averaged time series extracted from TC-following domains centered on IBTrACS best-track positions. Predictors are computed for multiple radial regions around the storm center to capture inner core and outer area conditions: inner-core variables are averaged over 0–2 (approximately 0–200 km), while outer-area predictors are averaged from 200–800 km or up to 1000 km, following conventions used in the extended SHIPS developmental dataset. To avoid backward-in-time causal dependencies, we only consider predictors at analysis time (00-hour), unlike operational SHIPS setup, which incorporates dynamical model forecast fields. As a result, forecast skill at longer lead times may be underestimated when synoptic-scale evolution is important. To expand the predictor set further, we use the Tropical Cyclone Precipitation, Infrared, Microwave, and Environmental Dataset (TC PRIMED) Razin and others (2023). TC PRIMED uses ERA5 to reconstruct operational SHIPS predictors as well as environmental and thermodynamic variables across multiple pressure levels and radial regions. An important limitation of the area-averaging approach is that it does not capture azimuthal asymmetries, which may be relevant for sheared or rapidly intensifying TCs Hazelton et al. (2017); Rogers et al. (2013); McNeely et al. (2020) and could be considered in future experiments.
@@
A key requirement for applying causal discovery to time series is the assumption of causal stationarity, i.e., the causal relationships between variables remain invariant over time within a given TC and across multiple storms in the multidata setting. This allows pooling of statistical evidence across time and the full training set for conditional independence testing Ganesh et al. (2023). Accordingly, TC time series are aligned relative to the time of minimum central pressure to satisfy the causal stationarity assumption. The mean sea level pressure (MSLP) time series for each storm in the training set are smoothed using a Gaussian filter ($\sigma=3\times 6\mathrm{\ hours}$) to reduce noise and the time of minimum central pressure is used as the reference point for alignment. For the shorter time series, we append NaN values on either sides to make sure that the lengths are consistent with the longest time series with minimum pressure in the middle so that the evolution of different storms could be compared on a common reference frame. All predictors are subsequently standardized using the mean and standard deviation computed from the training set. The final ERA5 dataset includes 214 predictors and the intensity-change target (DELV) at each forecast lead time. A total of 216 TCs from 2000–2019 are used within a storm-based cross-validation framework with predefined, non-overlapping training and validation splits, while an independent test set comprises 31 TCs from 2020–2021 including Hurricane Wilma (2005), retained as a particularly intense and challenging case.
@@ 2.3 SHIPS Developmental Dataset
To evaluate how causally informed predictors improve operational forecasting, we extract SHIPS predictors (Table S1) for the same 247 North Atlantic TC cases (2000-2021) used in the ERA5 experiments. We only used predictor values from the 00-hour initialization time in the SHIPS developmental dataset derived from GFS analysis fields, which represent the atmospheric conditions (i.e., observations assimilated into the GFS model) at forecast initialization time. This setup enables direct analogy between the SHIPS developmental dataset and the ERA5-derived dataset used for causal predictor discovery, as both rely on analysis time atmospheric fields within a purely statistical framework. We compare the generalization skill of statistical models trained using the original 21 SHIPS predictors (“original SHIPS”) with models trained on the same predictors augmented with causally selected predictors (“SHIPS+”). These additional predictors are identified through ERA5-based causal discovery experiments but are computed from GFS analysis fields in exactly the same way as the original SHIPS predictors. This setup tests whether the predictors selected in the ERA5 replication setting improve forecast skill when integrated into the operational SHIPS framework.
@@ 3.1 Causal Feature Selection
Causal feature selection is performed using the open-source TiGraMITe causal discovery framework Runge (2018); Runge et al. (2019b) which has been widely used to discover causal relationships from climate data Kretschmer et al. (2016, 2018); Samarasinghe et al. (2021). We employ the multidata PC (M-PC) algorithm, a modified version of the first step of the PCMCI algorithm Runge et al. (2019b); Runge (2020), to discover predictors of TC intensity change while eliminating spurious statistical associations. Unlike PCMCI+, which aims to infer the full causal graph including contemporaneous causal links, this work uses only the PC algorithm for causal feature selection Iglesias-Suarez et al. (2024); Runge et al. (2015). The “M” in M-PC indicates the multidata approach, where an ensemble of time series from multiple TCs is analyzed to identify a single set of causal predictors. Non-causal drivers within the ERA5 predictor set are detected and excluded through conditional independence test using linear partial correlation Ganesh et al. (2023). In our setup, this is done without considering time-lagged variables — each predictor is evaluated at a single timestep for the given forecast lead time. By filtering out non-causal predictors in the statistical models, this approach aims to improve the generalizability of statistical models to unseen TC cases.